OpenStudy (anonymous):

7. Write the equation y - 4 = -2(x - 5) in slope-intercept form.(Points : 1) y = -2x + 14 y = -2x + 5 y = -2x + 9 y = -2x + 6 8. If you are finding the line perpendicular to x + 3y = 5, what is the slope of the line?(Points : 1) -3/5 1/3 3 -3

OpenStudy (anonymous):

it always works the same way 1) distributive law \[ y - 4 = -2(x - 5)\] \[y-4=-2x+10\] 2) solve for \(y\) \[y=-2x+14\]

OpenStudy (anonymous):

the slope of \(ax+by=c\) is \(-\frac{a}{b}\) in your case the slope of \[x+3y=5\] is \(-\frac{1}{3}\) perpendicular line has slope that is the negative reciprocal

OpenStudy (anonymous):

negative reciprocal of \(-\frac{1}{3}\) is \(3\)